Papa’s Fried Chicken bought equipment on January 2, 2013, for $39,000. The equipment was expected to...

  1. Papa’s Fried Chicken bought equipment on January 2, 2013, for $39,000. The equipment was expected to remain in service for four years and to perform 11,000 fry jobs. At the end of the equipment’s useful life, Papa’s estimates that its residual value will be $6,000. The equipment performed 1,100 jobs the first year, 3,300 the second year, 4,400 the third, and 2,200 the fourth year.

Requirement: what is the depreciation and the book value per year for the equipment under the unit of production method and double-declining-balance method.

2.Assume that JR Tire Store completed the following perpetual inventory transactions for a line of tires:

May 1. Beginning merchandise inventory 16 tires @ $65 each

11 Purchase   10 tires @ $78 each

23 Sale 12 tires @ $90 each

26 Purchases 14 tires @ $80 each

29 sale 15 tires @ $90 each

Requirement:

  1. Compute cost of goods sold and gross profit using the FIFO inventory costing method.
  2. Compute cost of goods sold and gross profit using the LIFO inventory costing method.

Part III. Comprehensive Questions (15’*2=30’)

  1. During the whole December, Cruiser’s Boat Supply has the following transactions happened. (3’*6=18’)
  1. Dec 2, Cruiser’s Boat Supply Sold merchandise inventory to Walmart, $ 9200, on account. Terms 1/20, n/60. Cost of goods sold, $3500. FOB destination with the freight of $99, in cash.
  2. Dec 5, Cruiser’s Boat Supply purchased merchandise inventory with cash, $7900. FOB destination. Freight charged 75 of cash.
  3. Dec 16, Cruiser’s Boat Supply rent an office for its headquarter because of office renovation. The rent is paid already for $3,000 for three months.
  4. Dec 17, Cruiser’s Boat Supply returned $1200 of the merchandise purchased on December 5.
  5. Dec 21, Cruiser’s Boat Supply received the payment from Walmart.
  1. On Dec. 31, Cruiser’s Boat Supply made adjusting entries according to the following information. (3’*4=12’)
  1. In December, Cruiser’s Boat has delivered $5,000 of boats to its customer Costco, who prepaid 10,000 of cash in November. The cost of those boat, $2,000.
  2. The salaries of Cruiser’s Boat to its employees of December, $5,000 will be paid in next month.
  3. Depreciation of equipment for the month, $ 200
  4. Office rent which as prepaid on December 16 should be recorded for December, 2015.

Part IV. Conventional Question (10’)

  1. Explain the difference between FIFO and LIFO. Try to indicate the different consequences if a company uses them.
Answer #1

As per policy, only one question is allowed to answer at a time, but solving two questions for you here :

Q 1) Depreciation:
Under the unit of production method:
year rate (of 11000) Depreciation Book Value
0 39000
1 1100 =(39000-6000)*1100/11000=3300 35700
2 3300 9900 25800
3 4400 13200 12600
4 2200 6600 6000
Under the double declining balance method:
year rate Depreciation Book Value
0 39000
1 (100/4)*2=50% 19500 19500
2 (100/4)*2=50% 9750 9750
3 (100/4)*2=50% 3750 6000
4 (100/4)*2=50% 0 6000
Q2)
FIFO:
SALES = 12*90 + 15*90 =2430
COGS = 12*65+ (4*65 + 10*78 + 1*80)=1900
GROSS PROFIT = 2430 - 1900 = 530
LIFO:
SALES = 12*90 + 15*90 =2430
COGS = (10*78 + 2*65) + (14*80 + 1*10)=2040
GROSS PROFIT = 2430 - 2040 = 390